OpenStudy (anonymous):

how do u solve, log6(a^2)+2)+log2^2=2

OpenStudy (anonymous):

Can you rewrite ur question? it is not clear

OpenStudy (anonymous):

log 6^ (a^2)+2+log6^(2)=2

OpenStudy (anonymous):

log 6^[(a^2)+2]+log6^(2)=2

OpenStudy (anonymous):

oh that makes sense. Wait let me solve it

OpenStudy (anonymous):

Wait is the 6 meant to be subscript?

OpenStudy (anonymous):

Is six the base or is it a base of 10?

OpenStudy (anonymous):

Do you mean, solve \[\log_{6} (a ^{2}+2)+\log_{6}2=2\]?

OpenStudy (anonymous):

6 is the base, @lokisan yes!!:)

OpenStudy (anonymous):

So the answer is a = 4

OpenStudy (anonymous):

no i wrote that before and It came out incorrect

OpenStudy (anonymous):

6^2=2(a^2+2) 36=2(a^2+2) 18=a^2+2 16=a^2 4=a

OpenStudy (anonymous):

OK. \[\log_{6}(a^{2}+2)+\log_{6}2=\log_{6}[(a^{2}+2)(2)]=2\] so, by definition of the logarithm, \[(a^{2}+2)2=6^{2} {\rightarrow}a^{2}=16 {\rightarrow}a={\pm}4\]

OpenStudy (anonymous):

Oh ya. I forgot about that

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

You're welcome.